If 800ml of a juice drink is 10% grape juice, how much grape juice should be added to make a drink that is 20% grape juice?
Answer by stanbon(75887)

If 800ml of a juice drink is 10% grape juice, how much grape juice should be added to make a drink that is 20% grape juice?
----
Equation:
grape + graph = grape
0.10*800 + 1x = 0.20(800+x)
Multiply thru by 100 to get:
10*800 + 100x = 20*800 + 20x
80x = 10*800
x = 100 ml (amt. of pure grape needed)
=====================
Cheers,
Stan H.